OpenStudy (anonymous):

Simplify. square root of 3 times square root of 21

OpenStudy (owlcoffee):

Having this: \[\sqrt{3}\sqrt{21}\] I already proven the equality three times, so I will just write it: \[\sqrt[n]{a}\sqrt[n]{b}=\sqrt[n]{a.b}\] "a" and "b" are real numbers or variables, and "n" as well. So meaning that if I have the radical of any index multiplied by another radical of the same index but a different numebr inside, is the same as multiplying the numbers inside one whole radical. So, if we have: \[\sqrt{3}\sqrt{21}=\sqrt{3.21}\] having applied the property, now I'll let you multiply and finish the exercise.